# Latent Concept-based Explanation of NLP Models

### v1

@@ Abstract
Interpreting and understanding the predictions made by deep learning models poses a formidable challenge due to their inherently opaque nature. Many previous efforts aimed at explaining these predictions rely on input features, specifically, the words within NLP models. However, such explanations are often less informative due to the discrete nature of these words and their lack of contextual verbosity. To address this limitation, we introduce the Latent Concept Attribution method (LACOAT), which generates explanations for predictions based on latent concepts. Our foundational intuition is that a word can exhibit multiple facets, contingent upon the context in which it is used. Therefore, given a word in context, the latent space derived from our training process reflects a specific facet of that word. LACOAT functions by mapping the representations of salient input words into the training latent space, allowing it to provide predictions with context-based explanations within this latent space.
@@
Nadir Durrani Qatar Computing Research Institute, HBKU, Qatar ndurrani@hbku.edu.qa                        Hassan Sajjad Faculty of Computer Science, Dalhousie University, Canada hsajjad@dal.ca
@@ 1 Introduction
However, the explanation based solely on input words is less informative due to the discrete nature of words and the lack of contextual verbosity. A word consists of multifaceted aspects such as semantic, morphological, and syntactic roles in a sentence. Consider the word “trump” in Figure 1. It has several facets such as a verb, a verb with specific semantics, a named entity and a named entity representing a certain aspect such as tower names, family names, etc. We argue that given various contexts of a word in the training data, the model learns these diverse facets during training. Given a test instance, depending on the context a word appears, the model uses a particular facet of the input words in making the prediction. The explanation based on salient words alone does not reflect the facets of the word the model has used in the prediction and results in a less informed explanation.
@@
LACOAT is a local explanation method that provides an explanation of a single test instance. The reliance on the training data latent space makes the explanation reliable and further reflects on the quality of learning of the model and the training data. We perform qualitative and quantitative evaluation of LACOAT using the part-of-speech (POS) tagging and sentiment classification tasks across three pre-trained models. LACOAT generates an enriched explanation that is useful in understanding the model’s reasoning for a prediction. It also helps in understanding how the model has structured the knowledge of a task. We also conduct human evaluation to measure the utility of LACOAT with a human-in-the-loop.
@@ 2 Methodology
Consider a sentiment classification dataset and a sentiment classification model as an example. LACOAT works as follows: ConceptDiscoverer takes the training dataset and the model as input and outputs latent concepts of the model. At test time, given an input sentence, PredictionAttributor identifies the most salient input representations with respect to the prediction. ConceptMapper maps these salient input representations to the training data latent concepts and provides them as an explanation of the prediction. PlausiFyer takes the test sentence and its concept-based explanation and generates a human-friendly and insightful explanation of the prediction. In the following we describe these modules in detail.
@@ 2.1 ConceptDiscoverer
The goal of ConceptDiscoverer is to discover latent concepts given a model $\mathbb{M}$ and a dataset $\mathbb{D}$. We follow an identical procedure to Dalvi et al. (2022) to discover latent concepts. Specifically, for every word $w_{i}$ in $\mathbb{D}$, we extract contextual representations $\overrightarrow{z}_{w_{i}}$. We then cluster these representations using agglomerative hierarchical clustering Gowda and Krishna (1978). Specifically, the distance between any two representations is computed using the squared Euclidean distance, and Ward’s minimum variance criterion is used to minimize total within-cluster variance.
@@
2.2 Salient Representations Extraction
Given an input instance $s$, the goal of PredictionAttributor is to extract salient input representations with respect to the prediction $p$ from model $\mathbb{M}$. Gradient-based methods have been effectively used to compute the saliency of the input features for the given prediction, such as pure Gradient Simonyan et al. (2014), Input x Gradient Shrikumar et al. (2017) and Integrated Gradients (IG) Sundararajan et al. (2017b). In this work, we use IG as our gradient-based method as it’s a well-established method from literature. However, LACOAT is agnostic to the choice of the attribution method, and any other method that identifies salient input representations can be used while keeping the rest of the pipeline unchanged.
@@ 2.2 Salient Representations Extraction
In addition to the saliency based attribution method, we also experimented with using the position of the output head as an indication of the most salient contextual representation. A detailed description and results are provided in Appendix A.
@@ 2.4 PlausiFyer
Interpreting latent concepts can be challenging due to the need for diverse knowledge, including linguistic, task-specific, worldly, and geographical expertise (as seen in Figure 1). PlausiFyer offers a user-friendly summary and explanation of the latent concept and its relationship to the input instance using a Large Language Model (LLM).

### v2

@@ Abstract
Interpreting and understanding the predictions made by deep learning models poses a formidable challenge due to their inherently opaque nature. Many previous efforts aimed at explaining these predictions rely on input features, specifically, the words within NLP models. However, such explanations are often less informative due to the discrete nature of these words and their lack of contextual verbosity. To address this limitation, we introduce the Latent Concept Attribution method (LACOAT), which generates explanations for predictions based on latent concepts. Our foundational intuition is that a word can exhibit multiple facets, contingent upon the context in which it is used. Therefore, given a word in context, the latent space derived from our training process reflects a specific facet of that word. LACOAT functions by mapping the representations of salient input words into the training latent space, allowing it to provide latent context-based explanations of the prediction.
@@
WARNING: The appendix contains a few examples which may be disturbing to the reader
@@
Nadir Durrani Qatar Computing Research Institute, HBKU, Qatar ndurrani@hbku.edu.qa                        Marzia Nouri nouri.marzia.1999 @gmail.com                        Hassan Sajjad Faculty of Computer Science, Dalhousie University, Canada hsajjad@dal.ca
@@ 1 Introduction
However, the explanation based solely on input words is less informative due to the discrete nature of words and the lack of contextual verbosity. A word consists of multifaceted aspects such as semantic, morphological, and syntactic roles in a sentence. Consider the word “trump” in Figure 1. It has several facets such as a verb, a verb with specific semantics, and a named entity representing a certain aspect such as tower names, family names, etc. We argue that given various contexts of a word in the training data, the model learns these diverse facets during training. Given a test instance, depending on the context a word appears, the model uses a particular facet of the input words in making the prediction. The explanation based on salient words alone does not reflect the facets of the word the model has used in the prediction and results in a less informed explanation.
@@
LACOAT is a local explanation method that provides an explanation of a single test instance. The reliance on the training data latent space makes the explanation reliable and further reflects on the quality of learning of the model and the training data. We perform qualitative and quantitative evaluation of LACOAT using four classification tasks across four pre-trained models. LACOAT generates an enriched explanation that is useful in understanding the model’s reasoning for a prediction. We also conduct a human evaluation to measure the utility of LACOAT with a human-in-the-loop. Moreover, we measure the faithfulness of the most salient latent concept to the prediction using representation manipulation and show that it alters the prediction up to 46% of the time.
@@ 2 Methodology
Consider a sentiment classification dataset and a sentiment classification model as an example. LACOAT works as follows: ConceptDiscoverer takes the training dataset and the model as input and outputs latent concepts of the model. At test time, given an input sentence, PredictionAttributor identifies the most salient input representations with respect to the prediction. ConceptMapper maps these salient input representations to the training data latent concepts and provides them as an explanation of the prediction. PlausiFyer takes the test sentence and its concept-based explanation and generates a human-friendly and insightful explanation of the prediction.
@@ 2.1 ConceptDiscoverer
The goal of ConceptDiscoverer is to discover latent concepts given a model $\mathbb{M}$ and a dataset $\mathbb{D}$. We follow an identical procedure to Dalvi et al. (2022) to discover latent concepts. Specifically, for every word $w_{i}$ in $\mathbb{D}$, we extract contextual representations $\overrightarrow{z}_{w_{i}}$. We then cluster these representations using agglomerative hierarchical clustering Gowda and Krishna (1978). The distance between any two representations is computed using the squared Euclidean distance, and Ward’s minimum variance criterion is used to minimize total within-cluster variance.
@@
2.2 PredictionAttributor
Given an input instance $s$, the goal of PredictionAttributor is to extract salient input representations with respect to the prediction $p$ from model $\mathbb{M}$. Gradient-based methods have been effectively used to compute the saliency of the input features for the given prediction, such as pure Gradient Simonyan et al. (2014), Input x Gradient Shrikumar et al. (2017) and Integrated Gradients (IG) Sundararajan et al. (2017b). In this work, we use IG as our gradient-based method as it is a well-established method from literature. However, LACOAT is agnostic to the choice of the attribution method, and any other method that identifies salient input representations can be used while keeping the rest of the pipeline unchanged.
@@ 2.4 PlausiFyer
Interpreting latent concepts can be challenging due to the need for diverse knowledge, including linguistic, task-specific, worldly, and geographical expertise (as seen in Figure 1). PlausiFyer offers a user-friendly summary and explanation of the latent concept and its relationship to the input instance using a Large Language Model (LLM). Our intuition of natural language explanation is similar to Singh et al. (2023), however, we relied on latent concepts compared to most activated ngrams and the generation of synthetic data. Given an input sentence and the latent concept, we ask an LLM to explain the relationship between them. Due to space limitation, we present the prompts used for sequence labeling and classification tasks in App A.
3 Experimental Setup
Data
We use Parts-of-Speech (POS) Tagging, Toxicity classification (Toxicity), Sentiment Classification (Sentiment) and Natural Language Inference (NLI) tasks for our experiments. POS is a sequence labeling task while the other tasks are sequence classification tasks. We use the Penn TreeBank dataset Marcus et al. (1993) for POS, Jigsaw Toxicity dataset cjadams et al. (2017) for toxicity, the ERASER Movie Reviews dataset Pang and Lee (2004) for Sentiment and the MNLI dataset Wang et al. (2019) for the NLI tasks. Appendix B provides the information about each dataset.
Models
We fine-tune 12-layered pre-trained models; BERT-base-cased Devlin et al. (2019), RoBERTa-base Liu et al. (2019) and XLM-Roberta Conneau et al. (2020) using the training datasets of the tasks considered. For Llama2-2-7b-chat-hf Touvron et al. (2023), we use the base model without finetuning with zero-shot prompting for each task. We use transformers Wolf et al. (2020) with the default settings and hyperparameters. Task-wise performance of the models is provided in App. Tables 5,  613, and 17.
Module-specific hyperparameters
When extracting the activation and/or attribution of a word, we average the respective value over the word’s subword units. We optimize the number of clusters $K$ for each dataset as suggested by Dalvi et al. (2022). We use $K=600$ (POS, Toxicity) and $K=400$ (Sentiment, MNLI) for ConceptDiscoverer.
4 Evaluation
We perform a qualitative evaluation, a human evaluation and a module-level evaluation of LACOAT to measure its correctness and efficacy. We find consistent results across all tasks and models. Due to space limitation, we mainly present the results of POS and Sentiment using the BERT and RoBERTa models in the main paper. The full set of results are presented in Apps. HIJ.
4.1 Qualitative Evaluation
4.1.1 Evolution of Concepts
LACOAT generates the explanation for each layer with respect to a prediction. The layer-wise explanation shows the evolution of concepts in making the prediction. Figure 2 shows layers 0, 6 and 12’s latent concept of the most attributed input token for RoBERTa fine-tuned on the sentiment task (see App. Fig 5 for other examples). We found that the initial layer latent concepts do not always align with the sentiment of the input instance and may represent a general language concept. For instance, Figure 2(a) shows the concept of comparative and superlative adjectives of both positive and negative sentiments and is not limited to representing the negative sentiment of the most attributed word. In the middle layers, the latent concepts evolved into concepts that align better with the sentiment of the input sentence. For instance, the latent concept of Figure 2(b) shows a mix of adjectives and adverbs of negative sentiment, i.e. aligned with the sentiment of the input sentence. In the sentiment task, the most attributed word in the last layer is [CLS] which resulted in latent concepts consisting of [CLS] representations of the most related sentences to the input. In such cases, we randomly pick five [CLS] instances from the latent concept and show their corresponding sentences in the figure (see Figure 2(c)). We found that the last layer’s latent concepts are best aligned with the input instance and its prediction and are the most informative explanation of the prediction. In the rest of the paper, we focus our analysis on the explanations generated using the last layer only and perform a human evaluation to evaluate their efficacy and correctness.
4.1.2 Analyzing Last Layer Explanations
Correct prediction with correct gold label
Wrong prediction with correct gold label
To study how the explanation would change if it is a correct prediction, we employ TextAttack Morris et al. (2020) to create an adversarial example of the sentence in Figure 3(b) that flips its prediction. The new sentence replaces “laughing” with “kidding” which has a similar meaning but flipped the prediction to a correct prediction. Figure 6 in the App. shows the full explanation of the augmented sentence. With the correct prediction, the latent concept changed and the explanation clearly expresses a negative sentiment “… all express negative opinions and criticisms …" compared to the explanation of the wrongly predicted sentence.
Cross model analysis
LACOAT provides an opportunity to compare various models in terms of how they learned and structured the knowledge of a task. Figure 4 compares XLMR (top) and RoBERTa (bottom) for identical inputs. Both models predicted the correct label. However, their latent concept based explanation is substantially different. XLMR’s explanation shows a large and diverse concept where many words are related to finance and economics. RoBERTa’s latent concept is rather a small focused concept where the majority of tokens are units of measurement. It is worth noting that both models are fine-tuned on identical data.
Figure 4: Comparing explanation of XLMR (top) and RoBERTa (bottom)
4.2 Human Evaluation
We perform two human evaluations; one aimed at evaluating the usefulness of LACOAT’s explanation in understanding a prediction (LACOAT Effectiveness) and the other compares LACOAT with other explanation methods.
LACOAT Effectiveness
We conduct a human evaluation using four annotators across 50 test samples. Specifically, given an explanation (e.g. Figure 3), all annotators are asked to answer five questions (Q1-Q5) that aimed at evaluating the usefulness of LACOAT. [^1] We provide the evaluation questions in App. F. Specifically, Q1 evaluates whether LACOAT attributes the correct concept to a given prediction, while Q2 and Q3 measure the efficacy of LACOAT’s output in helping a user understand the prediction. Q4 and Q5 evaluates the output of PlausiFyer. They specifically separate out the cases where the explanation was accurate but irrelevant to the task at hand.
Comparison with other Methods
Despite a number of explanation methods proposed in the literature, it is hard to draw a comparison between them due to the difference in granularity of explanation, type of explanation and the methodology used. We design a human evaluation, asking evaluators to give a score between 1 to 3 to each of three explanations generated by IG, LACOAT and Cockatiel Jourdan et al. (2023). The annotation setup allows to rank multiple methods with the same usefulness rating. A total of 400 annotations were collected using four evaluators where each test instance is evaluated by all annotators. We provide the details of the evaluation setup and the results in App. F. The second part of Table 1 shows the percentage of samples for which each of the annotators ranked LACOAT to be the same or better than both IG and Cockatiel. The consolidated ranking is computed by averaging the ranks across users. The average Cohen’s $\kappa$ indicates Fair agreement between each annotator and the consolidated ranking. The results show that LACOAT explanation is more useful in understanding the prediction compared to other methods.
Top Labels Correct Samples Incorrect Samples All Samples
Annotation Fleiss $\kappa$
Q1 Yes/No 28 / 0 20 / 2 48 / 2 0.35
Q2 Helps/Neutral/Hinders 27 / 1 / 0 17 / 5 / 0 44 / 6 / 0 0.41
Q3 Helps/Neutral/Hinders 16 / 10 / 2 1 / 19 / 2 17 / 29 / 4 0.61
Q4 Yes/No 17 / 11 5 / 17 22 / 28 0.47
Q5 Yes/No 17 / 11 6 / 16 23 / 27 0.80
Bottom A1 A2 A3 A4 Consolidated Average Cohen’s $\kappa$
LACOAT $\uparrow$ 85% 72% 77% 87% 89% 0.37
Table 1: Top: Consolidated label distribution for Q1-Q5. Fleiss’ $\kappa$ scores are computed by averaging each annotator with the consolidated annotation. The consolidated labels and agreement scores are shown for all the samples, as well as partitioned into those where the model prediction was correct/incorrect. Bottom: Percentage of samples where LACOAT is ranked similar or better than other methods. A represents the average preference of LACOAT per annotator.
4.3 Module Specific Evaluation
POS Sentiment
Layers BERT RoBERTa BERT RoBERTa
9 92.38 86.97 31.94 99.59
10 92.79 89.64 99.57 99.69
11 93.39 89.95 99.71 99.48
12 93.95 90.04 99.25 99.27
Table 2: Accuracy of PredictionAttributor in mapping a representation to the correct latent concept.
4.3.1 ConceptDiscoverer
ConceptDiscoverer identifies latent concepts by clustering the representation. We question whether the discovered latent concepts are a true reflection of the properties that a representation possesses. Using ConceptDiscoverer, we form latent concepts of the last layer and automatically annotate them as described above. We found 87%, 83% and 86% of the latent concepts of BERT, RoBERTa and XLMR that perfectly map to a POS tag respectively. We further analyzed other concepts where 90% of the words did not belong to a single tag. We found them to be of compositional nature i.e. a concept consisting of related semantics like a mix of adjectives and proper nouns about countries such as Swedish and Sweden (App. Figure 9). For sentiment, we found 78%, 95%, and 94% of the latent concepts of BERT, RoBERTa, and XLMR to consist of either Positive or Negative sentences. The high number of class-based clusters of RoBERTa and XLMR show that at the 12th layer, the majority of their latent space is separated based on these two classes (see Table 7 for detailed results).
Layers 0 2 5 10 12
POS Top 1 100 100 99.03 92.67 84.19
Top 2 100 100 99.75 97.89 94.15
Top 5 100 100 99.94 99.68 99.05
Sentiment Top 1 100 100 97.19 83.09 68.24
Top 2 100 100 99.63 92.67 83.24
Top 5 100 100 99.94 97.75 94.24
4.3.2 PredictionAttributor
We question whether the salient input representation correctly represents the latent space of the output. This specifically evaluates PredictionAttributor. We calculate the number of times the representation of the most salient word/[CLS] token maps to the latent concept of the identical label as that of the prediction. We expect a high alignment at the top layers for PredictionAttributor to be correct. We do not include ConceptMapper when evaluating this and conduct the experiment using the training data only where we already know the alignment of a salient representation and the latent concept. Table 2 shows the results across the last four layers (See App. Tables 89 for full results). For POS, we observed a successful match of above 90% for all models. We observed the mismatched cases and found them to be also of a compositional nature i.e. latent concepts comprised of semantically related words (see App. Figure 9 for examples).
4.3.3 ConceptMapper
We evaluate the correctness of ConceptMapper in mapping a test representation to the training data latent concepts. ConceptMapper trains using representations and their cluster IDs as labels. We randomly split this training data into 90% train and 10% test where the test data serves as the gold annotation of latent concepts. We train ConceptMapper using the train instances and measure the accuracy of the test instances. Table 3 summarizes the results of BERT (See App. Tables 1011 for all results). Observing Top-1 accuracy, the performance of ConceptMapper starts high (100%) for lower layers and drops to 84.19 and 68.24% for the last layer. We found that the latent space becomes dense on the last layer. This is in line with Ethayarajh (2019) who showed that the representations of higher layers are highly anisotropic. This causes concepts to be close in the space. If true, the correct label should be among the top predictions of the mapper. We empirically tested it by considering the top two and top five predictions of the mapper, achieving a performance of up to 99.05% and 94.24% for POS and Sentiment respectively.
4.4 Faithfulness Evaluation
Zhao and Aletras (2023a) proposed masking parts of input token representations to evaluate faithfulness. We adapted their methodology to the latent concept faithfulness evaluation. We consider a salient latent concept highlighted by LACOAT to be faithful to the prediction if the ablation of that latent concept causes a change in prediction performance. We define ablation of a latent concept as removing the information of that latent concept from the prediction vector i.e. [CLS] . We calculate the vector of a latent concept by averaging the training representations that belong to the latent concept. At inference time, we subtract the latent concept vector from the [CLS] representation of layer 12 and perform the prediction. We report the accuracy of the model and the percentage of predictions altered (see Table 12 in App.). Moreover, we report the manipulation of [CLS] using random vectors. The results show a substantial change in all metrics when the latent concept is ablated compared to random, confirming the faithfulness of the latent concept based explanation.
5 Related work
The explainability methods can be approached by local explanations and global explanations targeting post-hoc analysis or introducing interpretability in the architecture Madsen et al. (2023); Sundararajan et al. (2017a); Denil et al. (2014); Selvaraju et al. (2020); Kapishnikov et al. (2021); Zhao and Aletras (2023b); Kim et al. (2018); Ghorbani et al. (2019); Jourdan et al. (2023); Zhao et al. (2023); Ribeiro et al. (2016); Rajagopal et al. (2021). Lyu et al. (2023) provides a survey of explainability methods in NLP. LACOAT is a local explanation method providing post-hoc explanations given an input instance. One of the common ways for local explanations is to interpret the model prediction based on the input features. However, such an explanation lacks contextual verbosity and it could not interpret the multifaceted roles of the input features.
Previous work attempted to explain and interpret NLP models using human-defined concepts Kim et al. (2018); Abraham et al. (2022) and concepts extracted from hidden representations Zhao et al. (2023); Ghorbani et al. (2019); Rajani et al. (2020); "Geva et al. (2022). Zhao et al. (2023); Kim et al. (2018) worked on the global explanation based on a surrogate model. We provide local explanations and we ensure the faithfulness of latent concepts by extracting them directly from the hidden representation without any supervised training. Rajani et al. (2020) used k-nearest neighbors of the training data to identify erroneous correlations and misclassified instances. Dalvi et al. (2022) analyzed latent concepts in their ability to represent linguistic knowledge. Our ConceptDiscoverer module is motivated by them. However, we propose a method to explain a model’s prediction using latent concepts.
6 Conclusion
We presented LACOAT that provides a faithful and human-friendly explanation of a model’s prediction. The qualitative evaluation and human evaluation showed that LACOAT explanations are insightful in explaining a correct prediction, in highlighting a wrong prediction and in comparing the explanations of models. The reliance on training data latent space enables interpreting how knowledge is structured in the network. Similarly, it enables the study of the evolution of predictions across layers. LACOAT promises human-in-the-loop in the decision-making process and is a step towards trust in AI.
7 Limitations
A few limitations of LACOAT are: 1) while hierarchical clustering is better than nearest neighbor in discovering latent concepts as established by Dalvi et al. (2022), it has computational limitations and it can not be easily extended to a corpus of say 1M tokens. However, the assumptions that are taken in the experimental setup e.g. considering the maximum 20 occurrences of a word (supported by Dalvi et al. (2022)) work well in practice in terms of limiting the number of tokens and covering all facets of a majority of the words. Moreover, the majority of the real-world tasks have limited task-specific data and LACOAT can effectively be applied in such cases. 2) For tasks requiring reasoning over multiple sentences, we observe that sometimes the LACOAT explanation’s are not clearly indicative of the reason of a prediction which might be based on some syntactic and semantic similarity between multiple input sentences. A possible solution to this is to consider hierarchical relationship between latent concepts in contrast to considering a flat structure among latent concepts. The underlying setup of ConceptDiscoverer supports this. However, comparing hierarchical structures requires further investigation beyond the scope of current work which provides a strong evidence towards faithful and human friendly explanations using training data latent space.
Appendix A Task-specific Prompts used with PlausiFyer